OpenStudy (anonymous):

-MEDAL WILL BE GIVEN- solve the equation for x [-x+4\ \left[\begin{matrix}2 & -8 \\ -4 & 2\end{matrix}\right]=\left[\begin{matrix}4 & -6 \\ 2 & -8\end{matrix}\right]

hero (hero):

Bro, you messed it up

OpenStudy (anonymous):

\[[-x+4\ \left[\begin{matrix}2 & -8 \\ -4 & 2\end{matrix}\right]=\left[\begin{matrix}4 & -6 \\ 2 & -8\end{matrix}\right]\\]\]

OpenStudy (anonymous):

fixed it